# Revisit What You See: Disclose Language Prior in Vision Tokens for Efficient Guided Decoding of LVLMs

### v1

@@ 3.1 Adaptive context-relevant vocabulary subset construction
At each decoding timestep $t$, LVLM $\mathcal{M}$ generate output distributions $p(h_{T+t-1}^{L},\mathcal{V})$ over the entire vocabulary space $\mathcal{V}$. However, as observed in Figure 2, the original distribution is often diffusely spread across many irrelevant tokens. Therefore, we dynamically constrain a vocabulary $\mathcal{V}$ into subset $\mathcal{V}_{\texttt{cons}}^{t}$ more suitable to capture the plausible semantic space of the output distribution. By restricting the output space to a smaller, contextually focused subset, this constraint reduce semantic distraction and facilitate more precise integration of visual information. To be specific, following [38], we adaptively define the vocabulary subset $\mathcal{V}_{\texttt{cons}}^{t}$ as below:
@@
$\mathcal{V}_{\texttt{cons}}^{t}=\left\{w\in\mathcal{V}:p(w|h_{T+t-1}^{L},\mathcal{V})\geq\alpha\cdot\max_{w^{\prime}}p(w^{\prime}|h_{T+t-1}^{L},\mathcal{V})\right\},$ (5)
@@
3.2 Vision token projection and selection
@@
$p(h_{i}^{j},\mathcal{V}_{\texttt{cons}}^{t}),\quad\text{for}\quad i\in\{0,\ldots,|\mathbf{v}|-1\},\;j\in\mathcal{J}.$ (6)
Intuitively, this distribution represents the likelihood of each candidate token within $\mathcal{V}_{\texttt{cons}}^{t}$ conditioned on the vision token embedding $h_{i}^{j}$. Next, to identify the vision token most relevant to the current decoding context, we compare each vision token distribution $p(h_{i}^{j},\mathcal{V}_{\texttt{cons}}^{t})$ against the vanilla output distribution $p(h_{T+t-1}^{L},\mathcal{V}_{\texttt{cons}}^{t})$ by computing the Jensen-Shannon Divergence (JSD):
$(i^{*},j^{*})=\arg\underset{i,j}{\min}\;\mathrm{JSD}\left(p(h_{T+t-1}^{L},\mathcal{V}_{\texttt{cons}}^{t})\parallel p(h_{i}^{j},\mathcal{V}_{\texttt{cons}}^{t})\right).$ (7)
@@
Remark. For the efficient implementation, the vision token projections over the full vocabulary $\mathcal{V}$, $p(h_{i}^{j},\mathcal{V})$ , are precomputed once before decoding begins. Then, at each timestep $t$, we apply slicing and masking to adapt the cached projections to the dynamically constrained subset $\mathcal{V}_{\texttt{cons}}^{t}$, enabling efficient selection without additional forward computation.
@@ 3.3 Output logit refinement using distribution from selected vision token
After identifying the most contextually relevant vision token at each decoding step, we then refine the vanilla output distribution guided by the reference visual grounding signal from the selected vision token. Specifically, given the vanilla output distribution $p(h_{T+t-1}^{L},\mathcal{V}_{\texttt{cons}}^{t})$ and the selected vision token distribution $p(h_{i^{*}}^{j^{*}},\mathcal{V}_{\texttt{cons}}^{t})$, we combine them through element-wise multiplication:
$p_{\tt{fin}}(y_{t})\propto\begin{cases}p_{y_{t}}(h_{T+t-1}^{L},\mathcal{V}_{\texttt{cons}}^{t})\times p_{y_{t}}(h_{i^{*}}^{j^{*}},\mathcal{V}_{\texttt{cons}}^{t}),&\text{if }y_{t}\in\mathcal{V}_{\texttt{cons}}^{t},\\ 0,&\text{otherwise}.\end{cases}$ (8)
@@ 4 Experiments
In this section, we design our experiments to investigate the following questions:
  • $\circ$
    Does ReVisiT improve visual grounding of LVLM during decoding? (Tables 1 and 2)
  • $\circ$
    Is ReVisiT generalizable to fine-grained hallucination evaluation? (Table 3)
  • $\circ$
    Is ReVisiT robust across various decoding methods and datasets? (Table 6)
  • $\circ$
    How much computational cost does ReVisiT save during inference? (Figure 3)
@@ 4.1 Setups
Datasets and evaluation metrics. To comprehensively assess the visual grounding capability of our method, we consider both generative and discriminative aspects of hallucination mitigation. Specifically, we evaluate our approach on three different benchmarks used to measure robustness and generalizability across vision-language tasks. 1) CHAIR [27] measures hallucination rates in image captioning by comparing generated captions to ground-truth object annotations from MS COCO [39] dataset. We report sentence-level (CHAIRS) and instance-level (CHAIRI) hallucination scores on 500 sampled images. In addition, we also report Recall, computed as the proportion of ground-truth objects correctly mentioned in the generated captions. 2) POPE [28] assesses visual grounding by posing binary object presence questions. We evaluate on MS COCO, A-OKVQA [40], and GQA [41] datasets constructing 3,000 samples per dataset and query type (random, popular, adversarial), with an equal split between positive and negative instances. We report accuracy, precision, and F1 score. 3) AMBER [29] provides fine-grained object-level datasets for both generative and discriminative evaluations of hallucination mitigation. We focus on the generative setting, reporting CHAIR, Cover, Hal, and Cog metrics over 1,004 samples to further analyze fine-grained . 4) LLaVA-Bench-In-the-Wild [6] consists of 24 diverse real-world images paired with 60 open-ended questions. We use this dataset with LLaVA-1.5-7B to qualitatively assess the ReVisiT’s robustness in complex, uncontrolled visual environments. Further dataset details and evaluation setups are provided in Appendix A.1.
@@
Models and baselines. To validate the generality of our method across different architectures and capabilities, we conduct experiments using two representative LVLMs: LLaVA-1.5-7B [12] as a widely adopted conventional baseline, and Qwen2.5-VL-7B [13] as a representative of state-of-the-art. As baselines, we adopt various intra-alignment decoding strategies to compare the effectiveness of our method in enhancing visual grounding. (a) Greedy decoding is a standard decoding baseline, and (b) DoLa [33] improves factuality by contrasting output logits from earlier and later decoder layers at each timestep, and has been widely adopted for hallucination mitigation in LVLMs despite being originally proposed for LLMs. For intra-alignment baselines, (c) VCD [23] mitigates hallucination by perturbing visual inputs with Gaussian noise and enforcing consistency between the original and perturbed outputs; (d) M3ID [24] contrasts outputs with and without visual input to alleviate language prior bias and enhance visual grounding; (e) SID [25] applies contrastive decoding on low-attention vision patches to mitigate hallucination and refine visual grounding.
@@
Implementation details. In all experiments, we apply greedy decoding (i.e., temperature is 0) with a maximum output length of 512 tokens for all decoding methods. Regarding baselines, we follow the default hyperparameters provided by their original implementations. For ReVisiT, we set the hyperparameters as follows: for LLaVA-1.5-7B, we select vision tokens mostly from the last decoder layer ("last"). For Qwen2.5-VL-7B, we select vision tokens from all even-numbered decoder layers ("all"). The APC threshold $\alpha$ is set for each task. More details are presented in Appendix A.2.
@@
Table 1: Results on CHAIR benchmark. Lower scores ($\downarrow$) on CHAIRS, CHAIRI and higher ($\uparrow$) recall indicate better performance. The best results in each setting are bolded, and the second-best are underlined.
Method LLaVA-1.5 Qwen2.5-VL
CHAIRS $\downarrow$ CHAIRI $\downarrow$ Recall $\uparrow$ CHAIRS $\downarrow$ CHAIRI $\downarrow$ Recall $\uparrow$
Greedy 53.8 14.66 80.81 32.2 8.51 $\underline{71.79}$
DoLA 53.6 14.45 80.23 35.2 8.14 69.80
VCD 52.8 15.61 80.17 37.8 9.33 $\underline{71.79}$
M3ID 57.0 16.57 81.96 32.8 9.40 68.78
SID 50.6 13.58 81.00 36.3 9.99 72.48
Ours 50.6 13.43 81.45 31.6 7.56 70.83
@@
Table 2: Average results on POPE benchmark. Average for 3 datasets and 3 types. Higher scores ($\uparrow$) on Accuracy, Precision, and F1 indicate better performance. The best results are bolded, and the second-best are underlined. Full results are presented in Table 6.
Method LLaVA-1.5 Qwen2.5-VL
@@
Greedy 79.47 74.48 82.36 84.49 93.34 82.82
DoLA 79.58 74.67 82.41 81.40 93.71 78.45
VCD 77.74 72.51 80.97 84.69 92.85 83.17
M3ID 79.48 74.48 82.37 83.93 92.32 82.31
SID 80.24 75.61 82.84 84.76 92.73 83.30
Ours 81.80 78.03 83.45 84.77 93.41 83.18
@@ 4.2 Main results
We first present the experimental results on CHAIR benchmark in Table 1. Notably, all baseline methods, including strong intra-alignment baselines (SID, M3ID, VCD) that explicitly attempt to ground visual semantics during decoding, underperform relative to vanilla greedy decoding on Qwen2.5-VL, highlighting the challenge of hallucination mitigation on stronger LVLMs. In contrast, ReVisiT consistently reduces the hallucination rates and achieves the best performance across both sentence-level (CHAIRS) and instance-level (CHAIRI) metrics for LLaVA-1.5 [12] and Qwen2.5-VL [13]. To be specific, compared to vanilla greedy decoding, ReVisiT reduces CHAIRS by 5.95% and CHAIRI by 8.39% on LLaVA-1.5, and achieves 1.86% and 11.16% reductions respectively on Qwen2.5-VL. Moreover, ReVisiT also attains competitive Recall scores, achieving the second-highest object coverage among all methods while effectively suppressing hallucinations.
Next, we present the results on POPE benchmark. Since POPE evaluates hallucination in a discriminative manner covering three datasets and three object types (random, popular, and adversarial), it poses a significant challenge for models to achieve robust and consistent performance across diverse settings. Table 2 summarizes the average accuracy across all 18 evaluation settings. Here, ReVisiT consistently achieves superior results, further confirming its effectiveness in mitigating hallucination across diverse object types and datasets. Particularly on GQA dataset under the popular and adversarial object settings, ReVisiT demonstrates its effectiveness even under more visually and semantically challenging scenarios, achieving absolute accuracy gains of over 5 percentage points compared to the second-best methods when using LLaVA-1.5. In addition, as shown in Table 6, ReVisiT outperforms vanilla greedy decoding across all 18 evaluation settings, demonstrating its strong robustness. In contrast, all existing baselines occasionally underperform vanilla greedy sampling which indicates achieving consistent improvements under POPE’s discriminative evaluation setting is non-trivial.
Furthermore, we evaluate ReVisiT and baseline approaches using the AMBER benchmark that provides a more recent and diverse evaluation of hallucination phenomena. Table 3 summarizes the evaluation results. Here, ReVisiT achieves an 8.33% reduction in the CHAIR score compared to vanilla greedy decoding on Qwen2.5-VL, while all other baselines exhibit higher CHAIR scores than the vanilla greedy. Given that CHAIR directly measures hallucination frequency and serves as the main evaluation criterion in the generative setting, we believe that it is the most important measurement for visual grounding quality. Across other metrics, ReVisiT demonstrates competitive or superior performance in general, which highlight the robustness and stability of our approach across diverse evaluation conditions. Overall, these results demonstrate that adaptively leveraging the most informative vision token at each decoding step, as proposed in ReVisiT, enables more faithful and visually grounded generation while maintaining simplicity.
Table 3: Results on AMBER benchmark. Lower scores ($\downarrow$) on CHAIR, Hal, Cog, and higher scores ($\uparrow$) on Cover indicate better performance. The best result for each metric is bolded, and the second-best is underlined.
Method LLaVA-1.5 Qwen2.5-VL
CHAIR $\downarrow$ Cover $\uparrow$ Hal $\downarrow$ Cog $\downarrow$ CHAIR $\downarrow$ Cover $\uparrow$ Hal $\downarrow$ Cog $\downarrow$
Greedy 7.2 50.6 32.7 3.8 4.8 64.0 27.8 1.6
DoLa 7.2 50.9 32.9 3.9 5.2 65.4 31.3 1.9
VCD 8.7 51.7 40.5 4.3 5.8 64.3 31.3 1.5
M3ID 7.2 57.0 40.7 3.0 4.9 53.4 21.8 0.9
SID 6.3 51.9 32.8 3.3 4.9 62.6 28.3 1.1
Ours 6.2 55.2 33.0 2.7 4.4 64.0 25.2 1.4
@@ 4.3 Additional analyses
In this section, we conduct additional analyses to further investigate the properties of ReVisiT.
@@
Inference speed improvement.
To evaluate the inference efficiency of ReVisiT compared to baseline decoding strategies, we measure the per-token computational time. All measurements are conducted on an NVIDIA RTX A6000 GPU with 48GB memory and an AMD EPYC 9354 32-Core Processor running at 3.25 GHz. To eliminate initial overhead, we perform 6 warmup iterations before measurement. During evaluation, each of the 300 samples is decoded three times, and the results are averaged to reduce variance. We report the average decoding time per generated token, measured in milliseconds (ms/token). Other evaluation configurations, including preprocessing pipelines and evaluation scripts, follow the official CHAIR benchmark settings. Exact numerical values are provided in the Appendix B.2.
As shown in Figure 3, ReVisiT achieves inference efficiency comparable to vanilla greedy decoding, exhibiting only 2.3% and 3.5% increases in computational time on LLaVA-1.5 and Qwen2.5-VL, respectively. Unlike ReVisiT, the intra-alignment baselines generally incur nearly 2×2$\times$2 × higher computational cost. In addition, DoLa, which requires an additional forward pass at every decoding step to project intermediate layers, also introduces 29.34% and 14.64% additional computation on LLaVA-1.5 and Qwen2.5-VL, respectively. These results highlight the key advantage of ReVisiT: it enhances visual grounding with negligible computational overhead, unlike prior methods that significantly compromise inference efficiency.

### v2

@@ 3.1 Adaptive context-aware vocabulary subset construction
At each decoding timestep $t$, LVLM $\mathcal{M}$ generate output distributions $p_{\theta}(h_{T+t-1}^{L},\mathcal{V})$ over the entire vocabulary space $\mathcal{V}$. However, as observed in Figure 2, the original distribution is often diffusely spread across many irrelevant tokens. Therefore, we dynamically constrain a vocabulary $\mathcal{V}$ into subset $\mathcal{V}_{\texttt{cons}}^{t}$ more suitable to capture the plausible semantic space of the output distribution. By restricting the output space to a smaller, contextually focused subset, this constraint reduce semantic distraction and facilitate more precise integration of visual information. To be specific, following Li et al. (2023a), we adaptively define the vocabulary subset $\mathcal{V}_{\texttt{cons}}^{t}$ as below:
@@
$\mathcal{V}_{\texttt{cons}}^{t}=\left\{w\in\mathcal{V}:p_{\theta}(w|h_{T+t-1}^{L},\mathcal{V})\geq\alpha\cdot\max_{w^{\prime}}p_{\theta}(w^{\prime}|h_{T+t-1}^{L},\mathcal{V})\right\},$ (3)
@@
3.2 Vision token projection and selection over subset
@@
$p_{\theta}(h_{i}^{j},\mathcal{V}_{\texttt{cons}}^{t}),\quad\text{for}\quad i\in\{0,\ldots,|\mathbf{v}|-1\},\;j\in\mathcal{J}.$ (4)
Intuitively, this distribution represents the likelihood of each candidate token within $\mathcal{V}_{\texttt{cons}}^{t}$ conditioned on the vision token embedding $h_{i}^{j}$. Next, to identify the vision token most relevant to the current decoding context, we compare each vision token distribution $p_{\theta}(h_{i}^{j},\mathcal{V}_{\texttt{cons}}^{t})$ against the vanilla output distribution $p_{\theta}(h_{T+t-1}^{L},\mathcal{V}_{\texttt{cons}}^{t})$ by computing the Jensen-Shannon Divergence (JSD):
$(i^{*},j^{*})=\arg\underset{i,j}{\min}\;\mathrm{JSD}\left(p_{\theta}(h_{T+t-1}^{L},\mathcal{V}_{\texttt{cons}}^{t})\parallel p_{\theta}(h_{i}^{j},\mathcal{V}_{\texttt{cons}}^{t})\right).$ (5)
@@
Remark. For the efficient implementation, the vision token projections over the full vocabulary $\mathcal{V}$, $p_{\theta}(h_{i}^{j},\mathcal{V})$ , are precomputed once before decoding begins. Then, at each timestep $t$, we apply slicing and masking to adapt the cached projections to the dynamically constrained subset $\mathcal{V}_{\texttt{cons}}^{t}$, enabling efficient selection without additional forward computation.
@@ 3.3 Output logit refinement using distribution from selected vision token
After identifying the most contextually relevant vision token at each decoding step, we then refine the vanilla output distribution guided by the reference visual grounding signal from the selected vision token. Specifically, given the vanilla output distribution $p_{\theta}(h_{T+t-1}^{L},\mathcal{V}_{\texttt{cons}}^{t})$ and the selected vision token distribution $p_{\theta}(h_{i^{*}}^{j^{*}},\mathcal{V}_{\texttt{cons}}^{t})$, we combine them through element-wise multiplication:
$p_{\tt{fin}}(y_{t})\propto\begin{cases}p_{\theta}(y_{t}|h_{T+t-1}^{L},\mathcal{V}_{\texttt{cons}}^{t})\times p_{\theta}(y_{t}|h_{i^{*}}^{j^{*}},\mathcal{V}_{\texttt{cons}}^{t}),&\text{if }y_{t}\in\mathcal{V}_{\texttt{cons}}^{t},\\ 0,&\text{otherwise}.\end{cases}$ (6)
@@ 4.1 Setups
Models and baselines. To validate the generality of our method across different architectures and capabilities, we conduct experiments using two representative LVLMs: LLaVA-1.5-7B (Liu et al., 2024a) as a widely adopted conventional baseline, and Qwen2.5-VL-7B (Bai et al., 2025) as a representative of state-of-the-art. As baselines, we adopt various intra-alignment decoding strategies to compare the effectiveness of our method in enhancing visual grounding. (a) Greedy decoding is a standard decoding baseline, and (b) DoLa (Chuang et al., 2024) improves factuality by contrasting output logits from earlier and later decoder layers at each timestep, and has been widely adopted for hallucination mitigation in LVLMs despite being originally proposed for LLMs. For intra-alignment baselines, (c) VCD (Leng et al., 2024) mitigates hallucination by perturbing visual inputs with Gaussian noise and enforcing consistency between the original and perturbed outputs; (d) M3ID (Favero et al., 2024) contrasts outputs with and without visual input to alleviate language prior bias and enhance visual grounding; (e) CODE (Kim et al., 2024) leverages the model’s self-generated image description as a contrasting reference to alleviate hallucination; (f) SID (Huo et al., 2025) applies contrastive decoding on low-attention vision patches to refine visual grounding.
@@
Implementation details. In entire experiments, we apply deterministic greedy decoding (i.e., $\text{temperature}=0$, $\text{top-}p=1$, $\text{top-}k=1$, $\text{num\_beams}=1$) with a maximum output length of 512 tokens for all decoding methods. Regarding baselines, we follow the default hyperparameters provided by their original implementations. For ReVisiT, we set the hyperparameters as follows: for LLaVA-1.5-7B, we select vision tokens mostly from the last decoder layer (”last”). For Qwen2.5-VL-7B, we select vision tokens mostly from all even-numbered decoder layers (”all”). The APC threshold $\alpha$ is set for each task. More details are presented in Appendix A.4.
@@
Table 2: HallusionBench. We report accuracies (%) on five metrics: qAcc (Question-pair accuracy), fAcc (Figure accuracy), Easy aAcc, Hard aAcc, and aAcc (All accuracy). Higher scores ($\uparrow$) indicate better performance for all metrics. he best results in each setting are bolded, and the second-best are underlined
Method LLaVA-1.5-7B Qwen2.5-VL-7B
qAcc $\uparrow$ fAcc $\uparrow$ Easy aAcc $\uparrow$ Hard aAcc $\uparrow$ aAcc $\uparrow$ qAcc $\uparrow$ fAcc $\uparrow$ Easy aAcc $\uparrow$ Hard aAcc $\uparrow$ aAcc $\uparrow$
Greedy 11.55 22.61 44.40 48.54 48.22 29.96 34.78 65.70 53.65 61.08
DoLA 13.36 22.61 45.85 49.27 49.24 27.08 30.87 62.09 51.46 58.04
VCD 14.44 22.61 44.77 47.45 47.55 29.60 35.22 64.62 54.74 61.25
M3ID 11.91 23.04 43.68 49.27 48.22 30.32 33.48 62.82 54.38 60.07
CODE 10.47 20.43 41.52 47.81 46.36 28.88 31.74 57.76 54.74 57.53
SID 12.27 21.30 44.77 48.18 48.22 32.13 36.09 65.70 55.11 61.59
[…]
@@
Table 3: CHAIR benchmark. Lower scores ($\downarrow$) on CHAIRS, CHAIRI and higher ($\uparrow$) F1 indicate better performance. The best results in each setting are bolded, and the second-best are underlined
Method LLaVA-1.5-7B Qwen2.5-VL-7B
CHAIRS $\downarrow$ CHAIRI $\downarrow$ F1 $\uparrow$ CHAIRS $\downarrow$ CHAIRI $\downarrow$ F1 $\uparrow$
Greedy 53.8 14.66 82.33 35.2 8.43 79.85
DoLA 53.6 14.45 82.27 31.0 7.63 79.33
VCD 52.8 15.61 81.84 35.0 9.07 80.10
M3ID 57.0 16.57 82.06 33.8 9.80 78.80
CODE 45.2 13.13 81.48 17.0 9.35 48.50
SID 50.6 13.58 83.34 37.6 8.84 80.40
[…]
Table 4: POPE benchmark. Average for 3 datasets and 3 types. Higher scores ($\uparrow$) on Accuracy, Precision, and F1 indicate better performance. The best results in each setting are bolded, and the second-best are underlined. Full results are presented in Table 11.
Method LLaVA-1.5-7B Qwen2.5-VL-7B
@@
Greedy 79.47 74.48 82.36 84.23 93.01 82.53
DoLA 79.58 74.67 82.41 81.36 92.36 78.71
VCD 77.74 72.51 80.97 84.43 92.48 82.91
M3ID 79.48 74.48 82.37 83.56 91.52 82.00
CODE 78.58 73.98 81.35 83.55 90.20 82.26
SID 80.40 75.94 82.91 85.78 92.27 84.72
[…]
Table 5: MMMU and VQAv2. For MMMU, we report Accuracy and Subject Win/Tie vs. Greedy; for VQAv2, we report Accuracy. Higher scores ($\uparrow$) indicate better performance for all metrics. The best results in each setting are bolded, and the second-best are underlined
Method MMMU VQAv2
LLaVA-1.5-7B Qwen2.5-VL-7B LLaVA-1.5-7B Qwen2.5-VL-7B
Accuracy $\uparrow$ Subj. Win/Tie Accuracy $\uparrow$ Subj. Win/Tie Accuracy $\uparrow$ Accuracy $\uparrow$
vs Greedy $\uparrow$ vs Greedy $\uparrow$
Greedy 34.39 - 51.11 - 69.47 65.80
DoLA 34.29 28 50.15 18 69.87 60.60
VCD 33.51 19 51.13 21 65.93 66.07
M3ID 34.51 29 46.63 15 70.33 59.87
CODE 34.48 22 42.79 5 66.00 46.60
SID 33.10 16 49.75 16 70.33 67.13
[…]
@@ 4.2 Main results
We evaluate both generative (caption faithfulness) and discriminative (object presence) grounding across five benchmarks to assess robustness and generalizability.
HallusionBench (Guan et al., 2024) probes image–context reasoning and entangled hallucinations with a LLM-based judge. We focus on the visual-dependent setting. Results are presented in Table 2, Across both LLaVA-1.5-7B and Qwen2.5-VL-7B, ReVisiT outperforms all baselines on the three representative metrics (qAcc, fAcc, and aAcc) among all methods. Notably, ReVisiT exceeds the second-best method by more than 5 points on qAcc, and reaches $1.75\times$ the score of vanilla greedy (from $11.55$ to $20.22$).
CHAIR (Rohrbach et al., 2018) measures hallucination rates in image captioning by comparing generated captions to ground-truth object annotations from MS COCO (Lin et al., 2014) dataset. We report sentence-level (CHAIRS) and instance-level (CHAIRI) hallucination scores on 500 sampled images, and additionally include F1 as an per-sample aggregate caption-level score. Table 3 shows that ReVisiT reduces hallucination while preserving or improving coverage on both models, maintaining at least second-best for all metrics.
POPE (Li et al., 2023b) evaluates object presence with binary questions across three datasets (MS COCO, A-OKVQA (Schwenk et al., 2022), and GQA (Hudson & Manning, 2019)) and three query types (random, popular, adversarial). We report Accuracy, Precision, and F1. As summarized in Table 4, ReVisiT ranks first or second in the overall averages across all datasets and query types, indicating robust gains in discriminative grounding.
MMMU (Yue et al., 2024) provides 30 subjects of VQA including both multiple-choice and open-ended settings. We report overall Accuracy and a per-subject win/tie vs. greedy rate to capture subject-wise gains over vanilla decoding. Results are presented in Table 5 (left). Across 30 subjects, ReVisiT attains the highest overall accuracy on both models while improving subject-wise robustness.
VQAv2 (Goyal et al., 2017) measures VQA correctness under inter-annotator variability via consensus accuracy over 10 annotations. As shown in Table 5 (right), ReVisiT improves general VQA performance and attains the best accuracy on both models: LLaVA-1.5-7B ($69.47\rightarrow\mathbf{70.40}$), Qwen2.5-VL-7B ($65.80\rightarrow\mathbf{67.60}$), surpassing the strongest baselines (e.g., $70.33$ and $67.13$).
Overall. ReVisiT succeeds across generative and discriminative settings, under both manually designed and LLM-aided evaluations, and on both hallucination-focused and general VQA tasks—without training, architectural changes, external modules, or multi-pass inference—by referencing intrinsic vision-token semantics. Comprehensive per-benchmark breakdowns, metrics, and extended tables are presented in Appendix A.2 and B.
@@ 4.3 Additional analyses
Inference speed improvement.
To evaluate the inference efficiency of ReVisiT compared to baseline decoding strategies, we measure the per-token computational time. All measurements are conducted on an NVIDIA RTX A6000 GPU with 48GB memory and an Intel(R) Xeon(R) Gold 5320 CPU running at 2.20 GHz. We perform 6 warmup iterations before measurement to eliminate initial overhead, and each of the 300 samples is decoded 3 times per sample during evaluation.
@@
We report the average decoding time per generated token, measured in milliseconds (ms/token). Other evaluation configurations, including preprocessing pipelines and evaluation scripts, follow the official CHAIR benchmark settings. Exact numerical values are provided in Table 12.
As shown in Figure 6, ReVisiT matches the inference efficiency of vanilla greedy decoding, increasing time by only $1.9\%$ on LLaVA-1.5-7B and $0.6\%$ on Qwen2.5-VL-7B. By contrast, methods that require additional forward passes (e.g., VCD / M3ID / SID / CODE) approach or exceed $2\times$ the greedy runtime. Even DoLa,which only adds a projection on intermediate layers each step, incurs $29.34\%$ and $14.64\%$ overhead on LLaVA-1.5-7B and Qwen2.5-VL-7B, respectively. These results highlight the key advantage of ReVisiT: it enhances visual grounding with negligible computational overhead, unlike prior methods that significantly compromise inference efficiency.
Figure 7: Ablation study. We evaluate the impact of vocabulary subset constraint, vision token selection, layer scope, and threshold $\alpha$ variation on CHAIR benchmark, reporting F1. Comprehensive results are presented in Table 10.